# AD-L-JEPA: Self-Supervised Spatial World Models with Joint Embedding Predictive Architecture for Autonomous Driving with LiDAR Data

### v1

@@ IV-C Downstream fine-tuning performance
Occupancy-MAE [1] 62.70 58.29 66.00 62.33 +0.23
AD-L-JEPA (ours) 62.68 58.39 66.48 62.51 +0.41
TABLE III: 3D detection results on the Waymo validation set with LEVEL_2 difficulty and the AP (%) metric. The models are pre-trained with and fine-tuned on the Waymo 20% data set.
Figure 4: Visual comparison of 3D object detection performance between SECOND trained from scratch and pre-trained with AD-L-JEPA, using 5%, 20%, and 100% of labeled data from the KITTI3D data set.
We first conduct self-supervised pre-training on KITTI3D without any labels and then fine-tune the network on the labeled training data from the same KITTI3D data set using the widely used SECOND [3] and PV-RCNN [33] methods. As shown quantitatively in Table I, our self-supervised pre-training approach significantly outperforms the baseline that corresponds to training from scratch and also surpasses other competitor algorithms. Table II presents ablation studies with different masking ratios during pre-training and their corresponding impact on downstream object detection performance for SECOND detection model and KITTI3D data set. We find that at a masking ratio of $0.5$, the difficulty is moderate and yields the best performance; thus, we set $0.5$ as the default masking ratio.
Next, in Table III we report the results of pre-training and fine-tuning with Waymo training data. The overall Average Precision across all classes outperforms the baseline trained from scratch, and our reproduced results using Occupancy-MAE. Notably, as the Waymo 20% data set is approximately $10$ times larger than KITTI, pre-training takes considerably longer, and there is a significant difference in the pre-training speed between AD-L-JEPA and Occupancy-MAE. Unlike Occupancy-MAE, which is a generative method involving computationally expensive dense $3$D convolutional layers to reconstruct invisible parts, AD-L-JEPA employs a joint-embedding predictive architecture at the BEV embedding level and does not require such layers. It takes about $12$ hours to pre-train using Occupancy-MAE for the default three epochs on $4$ $A100$ GPUs, totaling approximately $48$ $A100$ GPU hours. In contrast, pre-training with AD-L-JEPA for $30$ epochs on a single $A100$ GPU takes only $10$ hours, which translates to a significant improvement in computational efficiency.
IV-D Transfer learning and label efficiency performance
@@
TABLE IV: Transfer learning experiment: pre-training on Waymo ($20\%$ of the data set) and fine-tuning on KITTI3D. We report 3D detection results on the KITTI3D validation set with moderate difficulty, with the AP (%) metric. Label efficiency is studied when fine-tuning with $20\%$, $50\%$, and $100\%$ of the data. † denotes that the first layer of the model is randomly initialized.
@@
In this section, we report the transfer learning capability when moving from Waymo $20\%$ data set to KITTI3D with AD-L-JEPA.
@@
Each point in the Waymo $20\%$ data set features five attributes: [’$x$’, ’$y$’, ’$z$’, ’intensity’, ’elongation’], whereas each point in the KITTI3D data set includes four attributes: [’$x$’, ’$y$’, ’$z$’, ’intensity’]. Consequently, the first layer of the encoder pre-trained on the Waymo data set has a shape of $[5,16,3,3,3]$, while the first layer used for object detection on the KITTI3D data set has a shape of $[4,16,3,3,3]$ to accommodate the respective point features with a sparse $3$D convolutional kernel of $3$. The remaining layers have the same shape in both the pre-trained and downstream networks. When fine-tuning the downstream network with an encoder pre-trained directly on the Waymo $20\%$ data set without modifications, the first layer is randomly initialized, while the remaining layers of the $3$D encoder retain the pre-trained weights. In the bottom of Table IV, models denoted by † (Occupancy-MAE and AD-L-JEPA (ours)) present the quantitative results for this scenario using $100\%$ of the fine-tuning data. Surprisingly, even though the first layer is randomly initialized, the downstream task still benefits from positive transfers in case of AD-L-JEPA, while Occupancy-MAE experiences negative transfers. This experiment demonstrates that even if not all encoder layers are initialized from pre-trained weights, the intermediate pre-trained layers of our method are robust and continue to benefit the downstream tasks.

### v2

@@ Introduction
Occupancy-MAE, 20% 62.70 58.29 66.00 62.33 +0.23
Occupancy-MAE, 100% 63.04 58.81 66.55 62.80 +0.70
AD-L-JEPA (ours), 20% 62.68 58.39 66.48 62.51 +0.41
AD-L-JEPA (ours), 100% 63.07 58.64 66.81 62.84 +0.74
Table 2: 3D detection results on the Waymo validation set with LEVEL_2 difficulty defined in (Sun et al. 2020), the AP (%) and APH (%) metric. The models are pre-trained with $20\%$ or $100\%$ of the Waymo dataset, as indicated, and are fine-tuned on the Waymo 20% dataset.
Method Veh. Ped. Cycl. Overall Diff.
No pre-training 71.19 26.44 58.04 51.89
SECOND, pre-trained with 100k frames
BYOL 68.02 19.50 50.61 46.04 -5.85
PointContrast 71.07 22.52 56.36 49.98 -1.91
SwAV 72.71 25.13 58.05 51.96 +0.07
DeepCluster 73.19 24.00 58.99 52.06 +0.17
DepthContrast 71.88 23.57 56.63 50.69 -1.20
ProposalContrast 72.99 25.77 58.23 52.33 +0.44
ALSO 71.73 28.16 58.13 52.68 +0.79
Occupancy-MAE 73.54 25.93 58.34 52.60 +0.71
AD-L-JEPA (ours) 73.18 29.19 58.14 53.50 +1.61
SECOND, pre-trained with 500k frames
BYOL 70.93 25.86 55.63 50.82 -1.07
PointContrast 71.39 27.69 56.88 51.99 +0.10
SwAV 72.51 27.08 57.85 52.48 +0.59
DepthContrast 71.92 29.01 57.51 52.81 +0.92
DeepCluster 71.62 29.33 57.61 52.86 +0.97
AD-L-JEPA (ours) 73.25 31.91 59.47 54.87 +2.98
SECOND, pre-trained with 1M frames
BYOL 71.32 25.02 58.56 51.63 -0.26
PointContrast 71.87 28.03 60.88 53.59 +1.70
SwAV 72.46 29.84 57.84 53.38 +1.49
DeepCluster 72.89 30.32 57.94 53.72 +1.83
AD-L-JEPA (ours), 73.01 31.94 59.16 54.70 +2.81
Table 3: 3D detection results on the ONCE validation set with the AP (%) metric. General self‑supervised learning leads to negligle or even negative gains; by contrast, AD‑L‑JEPA pre‑trained with 500k frames significantly outperforms models pre‑trained with 1M frames.
Next, in Table 2, we report the results of pre-training with 20% or 100% of the Waymo training data set, and fine-tuning with 20% of the Waymo training data set with Centerpoint (Yin, Zhou, and Krahenbuhl 2021). The overall Average Precision across all classes outperforms the baseline trained from scratch, and our reproduced results using Occupancy-MAE.
Finally, Table 3 reports the results of pre‑training on the largest ONCE dataset followed by fine‑tuning with SECOND. General self-supervised learning methods yield negligible or even negative gains, indicating the need for a domain specific design. AD-L-JEPA pre-trained on 100k frames outperforms all models trained on 100k or 500k frames and nearly matches the best model trained on 1M frames, and when pre-trained on 500k frames it surpasses every competitor, demonstrating scalability and superior self-supervised learning capability. Interestingly, AD-L-JEPA pre-trained on 1M frames, although significantly better than other methods, falls slightly behind AD-L-JEPA pre-trained on 500K frames, demonstrating that it is a method that is efficient in terms of data and learns high level useful representations with less unlabeled data. This small drop aligns with existing literature showing that increasing the number of unlabeled samples consistently boosts performance but saturates at a point (Goyal et al. 2019). Such saturation can be explained by the data redundancy of highly similar driving scenarios in the 1M frame setting and could potentially be mitigated by increasing the diversity of pre-training data (Al Kader Hammoud et al. 2024).
Transfer Learning and Label Efficiency
@@
Table 4: Transfer learning experiment: pre-training on Waymo ($20\%$ of the data set) and fine-tuning on KITTI3D. We report 3D detection results on the KITTI3D validation set, with the AP (%) metric. Label efficiency is studied when fine-tuning with $20\%$, $50\%$, and $100\%$ of the data. † denotes that the first layer of the model is randomly initialized.
In this section, we report transfer learning from the Waymo 20% split to KITTI3D. Each Waymo point has five attributes, whereas each KITTI3D point has only four. When fine‑tuning with an encoder pre‑trained directly on the Waymo 20% split, we randomly initialize only the first layer (denoted at 100% labels) and retain the remaining 3D encoder weights. Alternatively, we drop the fifth ‘elongation’ feature, re‑pretrain on the first four attributes, and thus initialize every layer from pre‑trained weights. We also explore fine‑tune with 20%, 50%, and 100% of KITTI labels in this setting. Table 4 shows that AD‑L‑JEPA consistently outperforms baselines across different label efficiencies. Notably, $\text{AD-L-JEPA}^{\text{{\textdagger}}}$ model is strongest at 100%, likely because retaining pre‑training on all original five features yields richer representations.
@@
Other Evaluations
Occupancy Estimation
Singular Value Decomposition Analysis
We follow the methodology described in (Li, Efros, and Pathak 2022) to conduct singular value decomposition of non-empty BEV embeddings outputted by encoders pre-trained with AD-L-JEPA and Occupancy-MAE. This analysis assesses the level of dimensional collapse in pre-trained embeddings. Figure 5 shows that the normalized sorted singular values for AD-L-JEPA are more evenly distributed, while the corresponding cumulative explained variance increases more slowly, indicating less dimensional collapse and less redundant information in the embeddings.
Conclusions
In this paper, we propose AD-L-JEPA, the first joint-embedding predictive architecture for self-supervised representation learning of autonomous-driving LiDAR data. It learns useful representations without any labeled data by predicting masked regions in the BEV embedding space. It neither requires manual creation of positive/negative pairs for contrastive learning nor explicitly reconstructs the complex, high-uncertainty driving scenes. Extensive experiments show that AD-L-JEPA is a more efficient pre-training method for automotive LiDAR-based object detection than state-of-the-art methods, learning richer representations that generalize better to downstream tasks while requiring significantly fewer GPU pre-training hours and less GPU memory. For future work, we plan to extend AD-L-JEPA to leverage temporal dynamics and to incorporate action-conditioned self-supervised representation learning in autonomous driving scenarios.
Appendix
Appendix A Datasets Descriptions
We use the KITTI3D, Waymo, and ONCE datasets to conduct self-supervised pre‐training and downstream fine‐tuning.
KITTI3D contains 3,712 training frames and 3,769 validation frames. We report the 3D object detection Average Precision (AP) metric at 40 recall positions ($R_{40}$) and 11 recall positions ($R_{11}$) for the moderate difficulty level [^1] Moderate difficulty is defined in KITTI3D; LEVEL 2 difficulty, mentioned later, is defined in Waymo. across all classes on the KITTI3D dataset.
Waymo contains 158,081 training frames and 39,987 validation frames. Following the standard OpenPCDet protocol, we subsample 20% of the training frames at an interval of 5, yielding 31,617 frames. We report both the AP and the Average Precision with Heading (APH) metrics for LEVEL 2 difficulty across all classes on the Waymo dataset.
ONCE is a large‐scale unlabeled dataset designed for self‐supervised LiDAR pre‐training; it contains 1 million unlabeled frames. Depending on the subset, it offers small (100k), medium (500k), and large (1M) pools of unlabeled frames. In addition to this vast unlabeled corpus, it provides annotations for 16k LiDAR frames used in supervised downstream fine‐tuning.
Appendix B Experimental Details
The random seed is set to 666 for all experiments.
Pre-training Details
We use our proposed AD-L-JEPA to pre-train the popular VoxelBackBone8x encoder, which contains $12$ sparse convolution layers. For all dataset settings, we pre-train the network with unlabeled training data for $30$ epochs using the Adam optimizer and the default one-cycle learning rate scheduler in OpenPCDet, with a learning rate of $0.0003$ and a weight decay of $0.01$. We set the initial $\eta$ to 0.996 and increase it linearly to $1$, for the moving average update. We set masking ratio = $0.5$. We set $\alpha_{0}=0.25$ and $\alpha_{1}=0.75$ for empty grids and non-empty, respectively. We set $\beta_{1}=1$ and $\beta_{2}=1$. The threshold for variance regularization, $\gamma$, is set to $\frac{1}{\sqrt{D}}=\frac{1}{16}$ for $L_{2}$-normalized embeddings. For the KITTI3D dataset, we set $\lambda_{\text{jepa}}=1$ and $\lambda_{\text{reg}}=1$ to balance the two loss functions. For all settings of the Waymo dataset and ONCE dataset, we set $\lambda_{\text{jepa}}=1$ and $\lambda_{\text{reg}}=10$ to balance the two loss functions. We use four 1080 Ti GPUs for KITTI3D with a total batch size of 16, and a single A100 GPU with a batch size of 16 for Waymo during pre-training. For the largest ONCE dataset 100k setting, we use 4 A100 GPUs with a total batch size of 64, we use 8 A100 GPUs with a total batch size of 128 for 500k and 1M settings.
For pre-training experiments reproducing Occupancy-MAE, we follow all default settings in their paper for both pre-training and fine-tuning. We use 4 1080 Ti GPUs for the KITTI3D dataset, 4 A100 GPUs for the Waymo dataset, and 4 A100 GPUs for the ONCE dataset.
Fine-tuning Details
During fine-tuning, we initialize the encoder with the pre-trained context encoder’s weights and fine-tune the model for 80 epochs using the SECOND, PVRCNN, and CenterPoint detectors (all without residual connections), employing the default OpenPCDet hyperparameters. For each setup, we conduct three independent runs and report the best performance. Reproduced results may differ from those in the original papers. We use four 1080 Ti GPUs for KITTI3D and four RTX 4090 GPUs for Waymo (batch size 16). For the largest ONCE experiments, we use eight RTX 4090 GPUs (total batch size 64, i.e., 4 per GPU).
Figure 6: Visual comparison of 3D object detection performance between SECOND trained from scratch and pre-trained with AD-L-JEPA, using 20%, 50%, and 100% of labeled data from the KITTI3D data set.
Transfer Learning and Label Effiency Details
@@ Appendix B Experimental Details
For transfer learning experiments from Waymo 20% dataset to KITTI3D dataset: each point in the Waymo $20\%$ data set features five attributes: [’$x$’, ’$y$’, ’$z$’, ’intensity’, ’elongation’], whereas each point in the KITTI3D data set includes four attributes: [’$x$’, ’$y$’, ’$z$’, ’intensity’]. Consequently, the first layer of the encoder pre-trained on the Waymo data set has a shape of $[5,16,3,3,3]$, while the first layer used for object detection on the KITTI3D data set has a shape of $[4,16,3,3,3]$ to accommodate the respective point features with a sparse 33D convolutional kernel of $3$. The remaining layers have the same shape in both the pre-trained and downstream networks. When fine-tuning the downstream network with an encoder pre-trained directly on the Waymo $20\%$ data set without modifications, the first layer is randomly initialized, while the remaining layers of the 33D encoder retain the pre-trained weights.
Next, we also modify the first layer of the encoder used for pre-training on the Waymo $20\%$ data set to the shape of $[4,16,3,3,3]$ and redo the pre-training, utilizing only the first four features: ‘$x$’, ‘$y$’, ‘$z$’, and ‘intensity’, while ignoring the fifth feature ‘elongation’. This adjustment allows all layers of the encoder used for the downstream KITTI3D object detection task to be initialized from pre-trained weights. Furthermore, in this setting, for label efficiency of fine-tuning with varying amounts of labels, we report fine-tuning with $20\%$, $50\%$, and $100\%$ of the data and corresponding fine-tuning epochs of $320$, $160$, and $80$, respectively, to ensure a enough number of fine-tuning iterations.
Appendix C Additional Visualization
Figure 6 shows the qualitative comparisons with training from scratch using different amounts of labeled data for fine-tuning.
Appendix D Additional Ablation Studies
In terms of computation, we conduct all ablation studies on the KITTI3D dataset for both pre-training and fine-tuning, using the simple SECOND model for downstream object detection.
Ablation on Masking Ratios
Table 5 presents ablation studies with different masking ratios during pre-training and their corresponding impact on downstream object detection performance. We find that at a masking ratio of $0.5$, the difficulty is moderate and yields the best performance; thus, we set $0.5$ as the default masking ratio.
masking ratio 0.25 0.5 0.75
mAP - $R_{40}$ 66.21 67.92 67.24
mAP - $R_{11}$ 65.73 67.10 66.76
Table 5: Different masking ratio and the corresponding mAP (%) on KITTI3D data set with SECOND detection model.
Ablation on Learnable Empty/Mask Tokens and Explict Variance Regularization Loss
Please see the ablation studies in Table 6. Replacing representations of masked regions with a learnable mask token is a common practice in self-supervised learning. In LiDAR settings using a sparse convolutional network, empty regions (i.e., regions with no points) produce near-zero vectors after the target encoder, resulting in semantically similar embeddings. Without replacing these with a learnable empty token, direct $L_{2}$ normalization of such near-zero vectors yields high variance in the embeddings and causes them to diverge from their intended meanings. Additionally, when using explicit variance loss to prevent representation collapse, the downstream fine-tuning performance significantly improves. Given the above ablation studies, we introduce a learnable empty token, a learnable mask token, and explicit variance regularization as the default setting.
Learnable empty token
Learnable mask token
Variance regularization loss
mAP - $R_{40}$ 65.86 66.37 67.20 67.92
mAP - $R_{11}$ 65.48 65.91 66.77 67.10
Table 6: Ablation study of variance regularization loss, learnable empty token, and learnable mask token during pre-training, along with the corresponding fine-tuning mAP (%) on the KITTI3D dataset using the SECOND detection model.
Ablation on Predictor’s Depth
As the encoder’s architecture is fixed and thus can’t be changed for all settings, we conduct ablation studies on varying the predictor’s depth in Table 7 and observe empirically that depth=3 works the best. Therefore, we use that as the default setting for all experiments.
Number of layers 1 3 6
mAP - $R_{40}$ 67.58 67.92 67.17
mAP - $R_{11}$ 66.94 67.10 66.45
Table 7: Ablation study of predictor depth during pre-training, along with the corresponding fine-tuning mAP (%) on the KITTI3D dataset using the SECOND detection model.
Appendix E Detailed Pre-training Efficiency Comparisons
The detailed comparison of per-GPU memory consumption versus batch size between Occupancy-MAE and AD-L-JEPA can be found in Tables 8 and 9. AD-L-JEPA leads to $2.8\times$$3.4\times$ lower GPU memory usage for pre‑training on the 20% and 100% splits of the Waymo dataset, and $3.1\times$$4\times$ lower GPU memory usage for pre‑training on the ONCE 100k split.
Pre-training GPU Memory Comparison
Batch size 4 8 16
Occupancy-MAE 18.4 GB 34.6 GB 67.1 GB
AD-L-JEPA (ours) 6.7 GB 10.7 GB 19.6 GB
Table 8: Per-GPU memory (GB) consumption as batch size increases during pre-training on the Waymo dataset, measured on an A100 GPU.
Batch size 4 8 16
Occupancy-MAE 17.7 GB 33.7 GB 65.4 GB
AD-L-JEPA (ours) 5.7 GB 9 GB 16.5 GB
Table 9: Per-GPU memory (GB) consumption as batch size increases during pre-training on the ONCE dataset, measured on an A100 GPU.
Pre-training Time Comparison
Although, as mentioned in the previous section, AD-L-JEPA is efficient enough so that it can use only one A100 GPU for pre-training on the medium-scale Waymo dataset (whereas Occupancy-MAE requires four), we estimate the pre-training hours for both methods using 4 A100 GPUs on the Waymo dataset to enable a fair comparison. For the ONCE 100k dataset, we estimate the pre‐training hours for both methods using eight A100 GPUs. The detailed comparison of pre-training A100 hours consumption can be found in Tables 10. AD-L-JEPA reduces pre-training GPU hours by $2.7\times$ on the 20% and 100% Waymo splits and by $1.9\times$ on the ONCE 100k split.
Dataset Waymo 20% Waymo 100% ONCE 100k
Occupancy-MAE 56 hrs 210 hrs 56 hrs
AD-L-JEPA (ours) 19.25 hrs 77 hrs 30 hrs
Table 10: Pre‐training A100 GPU hours for the Waymo 20%, Waymo 100%, and ONCE 100k datasets.
Appendix F Pre-training logs
In this section, we present the pre-training logs over the course of training iterations for each setting.
For pre-training experiments on the KITTI3D dataset, each point has four features [’$x$’, ’$y$’, ’$z$’, ’intensity’], and the corresponding pre-training logs are shown in Figure 7; For pre-training experiments on the Waymo 20% and 100% datasets, each point has five features [’$x$’, ’$y$’, ’$z$’, ’intensity’, ’elongation’]. The logs for these experiments appear in Figure 8 and Figure 9, respectively.; Additionally, we conduct an experiment that uses only the first four point features [’$x$’, ’$y$’, ’$z$’, ’intensity’], excluding the fifth feature (elongation). Under this setting, the first layer of the pretrained encoder can be directly reused together with the other pretrained layers during fine-tuning. The logs for this experiment are shown in Figure 10.
In each setting, we track 12 values over the training iterations:
  • loss_pretrain: overall pre-training loss
  • loss_reg: overall variance regularization loss
  • loss_reg_prediction_target_voxels: variance regularization loss for predicted target BEV embeddings at non-empty regions
  • loss_reg_context_context_voxels: variance regularization loss for context BEV embeddings at non-empty regions
  • loss_jepa: overall joint-embedding prediction loss
  • loss_cos_jepa_target_voxels: joint-embedding prediction loss for non-empty regions
  • loss_cos_jepa_target_empty_voxels: joint-embedding prediction loss for empty regions
  • var_target_target_voxels: variance for non-empty regions in target BEV embeddings
  • var_prediction_target_voxels: variance for non-empty regions in predicted target BEV embeddings
  • var_prediction_target_empty_voxels: variance for empty regions in predicted target BEV embeddings
  • var_context_context_voxels: variance for non-empty regions in context BEV embeddings
  • learning_rate: learning rate
Figure 7: Pretraining logs using AD-L-JEPA on the KITTI3D dataset (best viewed when zoomed in).
Figure 8: Pretraining logs using AD-L-JEPA on the Waymo 20% dataset (best viewed when zoomed in).
Figure 9: Pretraining logs using AD-L-JEPA on the Waymo 100% dataset (best viewed when zoomed in).